OpenStudy (anonymous):

Which of the following circles have their centers on the x-axis? Check all that apply. A. (x - 2)2 + (y + 3)2 = 2 B. (x - 0)2 + (y - 3)2 = 23 C. (x - 0)2 + (y - 0)2 = 72 D. (x - 4)2 + (y - 0)2 = 74

OpenStudy (anonymous):

Not A

OpenStudy (anonymous):

or D

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

It is D

OpenStudy (anonymous):

so what is t

OpenStudy (anonymous):

it is D

OpenStudy (anonymous):

only D

OpenStudy (ybarrap):

This, \((x -h)^2+(y-k)^2=r^2 \), indicates that the circle is centered at\(<h,k>\). For each options above, find where \(k=0\), that is, the y coordinate of the center of the circle does not shift above or below the x-axis and those are your answers.

OpenStudy (amistre64):

given a center (a,b) a circle is defined to as: \[(x-a)^2+(y-b)^2=r^2\] since the x axis is the y=0 axis .... \[(x-a)^2+(y-0)^2=r^2\] will be centered someplace along the x axis

OpenStudy (anonymous):

so C and D

OpenStudy (ybarrap):

yes